# Ramsey goodness of stars and fans for the Hajós graph

### v1

@@
Abstract: Given two graphs $G_{1}$ and $G_{2}$, the Ramsey number $R(G_{1},G_{2})$ denotes the smallest integer $N$ such that any red-blue coloring of the edges of $K_{N}$ contains either a red $G_{1}$ or a blue $G_{2}$. Let $G_{1}$ be a graph with chromatic number $\chi$ and chromatic surplus $s$, and let $G_{2}$ be a connected graph with $n$ vertices. The graph $G_{2}$ is said to be Ramsey-good for the graph $G_{1}$ (or simply $G_{1}$-good) if, for $n\geq s$,
@@
Keywords: Ramsey goodness, the Hajós graph, fan
AMS Subject Classification: 05C55, 05D10
@@ 1 Introduction
The study of Ramsey numbers for sparse graphs has flourished since the 1970s. At that time, Burr [2] established a general lower bound for the Ramsey number of any pair of graphs. Suppose $G_{2}$ is a connected graph, and let $\chi(G_{1})$ and $s(G_{1})$ denote the chromatic number and the chromatic surplus of $G_{1}$, respectively, where the chromatic surplus refers to the smallest size of a color class over all proper $\chi(G_{1})$-colorings of $G_{1}$. Then, if $|G_{2}|\geq s(G_{1})$,
@@
When equality holds in (1), the graph $G_{2}$ is said to be $G_{1}$-good, or equivalently, $G_{2}$ is Ramsey-good for $G_{1}$. In the special case where $G_{1}$ is the complete graph $K_{k}$, such graphs $G_{2}$ are referred to as $k$-good by Burr and Erdős [3].
When $G_{1}$ is the complete $(k+1)$-partite graph $K_{1,m_{1},m_{2},\ldots,m_{k}}$ and $G_{2}$ is a star, Burr, Faudree, Rousseau, and Schelp [4] proved that for $m=\min\{m_{i}\mid i\in[k]\}$ and sufficiently large $n$,
@@
A frequently studied class of star-like graphs is the fan. A fan $F_{n}$ is the graph formed by $n$ triangles sharing a common vertex; see Figure 1. The concept of the fan was first introduced into extremal graph theory by Erdős, Füredi, Gould, and Gunderson [9], who studied its extremal graph and Turán number. For results on the Ramsey numbers of fans, see [5, 8, 15, 21].
@@
Li and Rousseau [14] investigated the Ramsey-goodness of generalized fans. For any graphs $H$ and $G$, they proved that $K_{1}+nH$ is Ramsey-good for $K_{2}+G$ when $n$ is sufficiently large.
Observe that in any proper coloring of the graph $K_{2}+G$, there are two color classes containing only one vertex each. What happens if we require only one color class to have a single vertex? Let $K_{1}+K_{k}(m)$ denote the complete $(k+1)$-partite graph in which one partite set has size $1$ and each of the remaining $k$ partite sets has size $m$. Lin, Li, and Dong [16] showed that for $k\geq 2$, if $m$ is odd or if $n|H|$ is odd, then
@@
Subsequently, without relying on the Erdős-Stone-Simonovits stability theorem, Chung and Lin [7] determined the exact value of $R(K_{1,m_{1},m_{2},\ldots,m_{k}},K_{1}+nH)$ for large $n$ in all cases. Let $m=\min\{m_{i}\mid i\in[k]\}$. Then
@@
Another frequently studied star-like graph is the book $B_{k,n}$, which is the graph $K_{k}+(n-k)K_{1}$. Equivalently, $B_{k,n}$ can be viewed as the graph obtained by blowing up the center of the star $K_{1,n-k}$ into a clique $K_{k}$, while preserving the adjacency between each vertex of $K_{k}$ and the remaining vertices. For results concerning the Ramsey goodness of $B_{k,n}$, we refer the reader to [12, 19, 18]. For results on the Ramsey non-goodness of $B_{k,n}$, see [10, 11, 17].
@@
At the end of this section, we introduce some concepts and notation that will be used throughout. We use $[n]$ to denote the set $\{1,2,\ldots,n\}$. For a graph $G$ and a vertex subset $X\subseteq V(G)$, the subgraph induced by $X$ is denoted by $G[X]$, whose vertex set is $X$ and whose edge set consists of all edges in $G$ with both endpoints in $X$. For $u\in V(G)$ and $X\subseteq V(G)$, let $N_{X}(u)$ denote the set of neighbors of $u$ in $X$, and let $|N_{X}(u)|$ denote the number of such neighbors, also written as $d_{X}(u)$. When $X=V(G)$, we write $N_{X}(u)$ and $d_{X}(u)$ simply as $N(u)$ and $d(u)$, respectively. The minimum and maximum degrees of $G$ are denoted by $\delta(G)$ and $\Delta(G)$, respectively. Notation and terminology not explicitly defined follow Bondy and Murty [1].
@@ 2 Proof of Theorem 1
When $n$ is odd, let $\ell=(n+1)/2$. Consider the graph $(\ell K_{2})+(\ell K_{2})$, which consists of two disjoint copies of a matching with $\ell$ edges, where the edges between the two copies form a complete bipartite graph. It is straightforward to verify that this graph does not contain the Hajós graph as a subgraph, and that the maximum degree in its complement is $n-1$. Therefore, $R(H_{a},K_{1,n})\geq 2n+3$ for odd $n$.
@@
By the classical result of Chvátal [6], we have $R(K_{3},K_{1,n})=2n+1$. Therefore, the graph $G$ must contain a triangle; without loss of generality, let its vertices be $u_{1},u_{2},u_{3}$. For $1\leq i<j\leq 3$, the minimum degree condition implies
@@
We first consider the case when $n$ is even. Since $\delta(G)\geq n+2$, each vertex $w_{i}$ for $i\in[4]$ has at least $n-1$ neighbors in $V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}$. As $|V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}|=2n-2$, there must exist two vertices, say $w_{1}$ and $w_{2}$, that share a common neighbor in $V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}$, denoted by $w_{5}$. Since $G$ does not contain $K_{5}-e$ as a subgraph, neither $w_{3}$ nor $w_{4}$ is adjacent to $w_{5}$.
Let $X=V(G)\setminus\{w_{1},w_{2},w_{3},w_{4},w_{5}\}$. Then $|X|=2n-3$. By the minimum degree condition,
@@
Since $|X|=2n-3$, $d_{X}(w_{1})\geq n-2$ and $d_{X}(w_{3})\geq n-1$, the sets $N_{X}(w_{1})$ and $N_{X}(w_{3})$ form a partition of $X$, and we have $|N_{X}(w_{1})|=n-2$ and $|N_{X}(w_{3})|=n-1$. Given that $d_{X}(w_{2})\geq n-2$ and $N_{X}(w_{2})\cap N_{X}(w_{3})=\emptyset$, it follows that $N_{X}(w_{1})=N_{X}(w_{2})$. Therefore,
@@
Since $n$ is even, by the classical result of Harary [13],
@@
Now consider the case where $n$ is odd. Let $Y=V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}$. Since $\delta(G)\geq n+3$, each vertex $w_{i}$ for $i\in[4]$ has at least $n$ neighbors in $Y$. Given that $|Y|=2n-1$, any two vertices among $w_{1},w_{2},w_{3},w_{4}$ share at least one common neighbor in $Y$. Without loss of generality, assume that $w_{1}$ and $w_{2}$ share a common neighbor $w_{5}$ in $Y$, and $w_{2}$ and $w_{3}$ share a common neighbor $w_{6}$ in $Y$. As $G$ does not contain $K_{5}-e$ as a subgraph, vertices $w_{5}$ and $w_{6}$ must be distinct. Consequently, we obtain a Hajós graph in which the vertex subsets $\{w_{1},w_{2},w_{3}\}$, $\{w_{1},w_{2},w_{5}\}$, $\{w_{2},w_{3},w_{6}\}$, and $\{w_{1},w_{3},w_{4}\}$ each induce a triangle. ∎
3 Proof Theorem 2
The proof of Theorem 2 relies on the following key lemma, which was established by Zhao and the latter two authors of this paper [20].
@@ 3 Proof Theorem 2
To prove the upper bound, we proceed by contradiction. Assume that a graph $G$ with $24n+$ 2 vertices contains neither $H_{a}$ as a subgraph nor $F_{n}$ as a subgraph in $\overline{G}$. We will derive contradictions in the following two cases.
@@
Let $u$ be a vertex in $\overline{G}$ with the maximum degree. Denote by $H^{{}^{\prime}}$ the subgraph of $\overline{G}$ induced by the set of vertices adjacent to $u$ in $\overline{G}$, i.e., $H^{{}^{\prime}}=\overline{G}[N_{\overline{G}}(u)]$. Let $M$ be a maximum matching in $H^{{}^{\prime}}$. Then $M$ must contain exactly $n-1$ edges. This is because if $M$ has at least $n$ edges, these $n$ edges together with $u$ would form an $F_{n}$ in $\overline{G}$, a contradiction. Conversely, if $M$ has at most $n-2$ edges, the number of vertices in $H^{{}^{\prime}}$ not covered by $M$ (i.e., the remaining vertices after removing the vertices of $M$) is at least $62n+2-2(n-2)=$ 6. These vertices form a complete graph in $G$, which clearly contains $H_{a}$ as a subgraph, leading to a contradiction.
@@
We only need to prove that $u_{1}u_{3}\not\in E(G)$; the other case can be proven similarly. Suppose $u_{1}u_{3}\in E(G)$. Since $\delta(G)\geq 2n$, the set $\{u_{1},u_{2},u_{3}\}$ must contain two vertices, say $u_{1}$ and $u_{3}$, that have a common neighbor in $V(G)\setminus V(W_{4})$. Otherwise, this would contradict the fact that $G$ has $24n+$ 2 vertices. Let $u_{5}$ be a common neighbor of $u_{1}$ and $u_{3}$ in $V(G)\setminus V(W_{4})$. Then the subgraph induced by the vertex set $\{u_{0},u_{1},u_{2},u_{3},u_{4},u_{5}\}$ in $G$ contains $H_{a}$ as a subgraph, a contradiction. ∎
If $u_{1}$ and $u_{2}$ have a common neighbor distinct from $u_{0}$, denoted by $u_{5}$, then the subgraph induced by the vertex set $\{u_{0},u_{1},u_{2},u_{3},u_{4},u_{5}\}$ contains $H_{a}$ as a subgraph. Therefore, the only common neighbor of $u_{1}$ and $u_{2}$ can be $u_{0}$. Furthermore, since $\delta(G)\geq 2n$, $u_{2}$ must have at least $12n-$ 1 neighbors in $V(G)\setminus N(u_{1})$. Similarly, $u_{4}$ must also have at least $12n-$ 1 neighbors in $V(G)\setminus N(u_{1})$. Hence, the number of common neighbors of $u_{2}$ and $u_{4}$ in $V(G)\setminus N(u_{1})$ is at least
$42(2n-1)-|V(G)\setminus N(u_{1})|\geq 2(2n-1)-(4n+2-2n)=2n-$ 4 .
@@
Next, we only need to prove that the graph $G[U_{1}\setminus\{u_{1},u_{3}\}]$ does not contain a triangle as a subgraph. Again, we proceed by contradiction. Suppose $G[U_{1}\setminus\{u_{1},u_{3}\}]$ contains a triangle with vertices $x_{1},x_{2},x_{3}$. There must exist two vertices among $\{x_{1},x_{2},x_{3}\}$, say $x_{1}$ and $x_{3}$, that have a common neighbor in $V(G)\setminus\{x_{1},x_{2},x_{3},u_{2},u_{4}\}$, otherwise this would contradict the fact that the number of vertices in $G$ is $24n+$ 2. Let $x_{4}$ be a common neighbor of $x_{1}$ and $x_{3}$ in $V(G)\setminus\{x_{1},x_{2},x_{3},u_{2},u_{4}\}$. Then the subgraph induced by the vertex set $\{x_{1},x_{2},x_{3},x_{4},u_{2},u_{4}\}$ in $G$ contains $H_{a}$ as a subgraph, which leads to a contradiction. ∎
@@
When $k\geq 2$, for each $11\leq i\leq k-$ 1, the center vertex of $C_{i}$ and a leaf vertex of $C_{i+1}$ form an edge in $\overline{G}$. Similarly, the center vertex of $C_{k}$ and a leaf vertex of $C_{1}$ form an edge in $\overline{G}$. These $k$ edges form a matching in $\overline{G}$. The remaining vertices in $\overline{G}[U_{1}\setminus\{u_{1}\}]$ induce a complete subgraph, allowing us to find a matching of $n$ edges in $\overline{G}[U_{1}\setminus\{u_{1}\}]$. This gives a fan $F_{n}$ centered at $u_{1}$ in $\overline{G}$, a contradiction.
@@
Next, consider $w_{1},w_{2}\in W_{3}$. By the definition of $W_{3}$, $w_{1}$ and $w_{2}$ each have at least $92n-$ 9 neighbors in $U_{2}$. Thus, $w_{1}$ and $w_{2}$ share at least $182(2n-9)-|U_{2}|\geq 2n-$ 18 common neighbors in $U_{2}$, where this inequality follows from Claim 3.
If each common neighbor of $w_{1}$ and $w_{2}$ in $U_{2}$ has degree at least $2$ in $G[U_{2}]$, note that each connected component in $G[U_{2}]$ has at most one vertex of degree at least $2$. Hence, $G[U_{2}]$ must have at least $182n-$ 18 components, each containing a vertex of degree at least $2$, contradicting $|U_{2}|\leq 2n$. Thus, there exists a common neighbor $u_{z}$ of $w_{1}$ and $w_{2}$ in $U_{2}$ with degree at most $1$ in $G[U_{2}]$. Then the vertex $u_{z}$ has at least $112n-2-|W|\geq 2n-$ 11 neighbors in $U_{1}$. Consequently, $u_{z}$, $w_{1}$, and $w_{2}$ share at least $(2n-11)+(|U_{1}|-5)-|U_{1}|\geq 2n-16$ common neighbors in $U_{1}$. Selecting three such common neighbors together with $u_{z},w_{1},w_{2}$ induces a complete multipartite graph $K_{1,1,1,3}$, which contains $H_{a}$ as a subgraph, a contradiction.
If $w_{1}$ and $w_{2}$ belong to $W_{1}$ and $W_{3}$, respectively, let $w_{1}\in W_{1}$ and $w_{2}\in W_{3}$. Then $w_{1}$ must be adjacent to either $u_{2}$ or $u_{4}$. By symmetry, assume $w_{1}u_{2}\in E(G)$. According to the selection of $W_{3}$, $w_{2}$ has at least $92n-$ 9 neighbors in $U_{2}$. Therefore, $w_{2}$ must have a neighbor $u_{z}$ in $U_{2}$ whose degree in $G[U_{2}]$ is at most 1; otherwise, this would contradict $|U_{2}|\leq 2n$. Consequently, the number of neighbors of $u_{z}$ in $U_{1}$ is at least $112n-2-|W|\geq 2n-$ 11. Thus, the number of common neighbors of $u_{z}$, $w_{1}$, and $w_{2}$ in $U_{1}$ is at least $(2n-11)+(|U_{1}|-5)-|U_{1}|\geq 2n-16$. Selecting two common neighbors, together with $u_{z}$, $w_{1}$, $w_{2}$, and $u_{2}$, forms a subgraph in $G$ that contains $H_{a}$ as a subgraph, leading to a contradiction. ∎
@@
If exactly one vertex in $\{w_{1},\ldots,w_{t}\}$ is not in $V(M_{1})$, assume without loss of generality that $w_{1}\notin V(M_{1})$. By the choice of $M_{1}$, $w_{1}$ is adjacent to all vertices in $U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))$, including $v_{0}$, $v_{1}$, and $v_{2}$. This implies that $w_{1}$ has at least $112n-$ 11 neighbors in $U_{1}$.
If $w_{1}\in W_{3}$, by the definition of $W_{3}$, $w_{1}$ has at least $112n-$ 11 neighbors in $U_{2}$. Since $v_{2}$ has exactly one neighbor in $U_{1}$, it must have at least $112n-2-|W|\geq 2n-$ 11 neighbors in $U_{2}$. Thus, $w_{1}$ and $v_{2}$ share at least 2($222(2n-11)-|U_{2}|\geq 2n-$ 22 common neighbors in $U_{2}$. Let $u^{\prime}$ be one such common neighbor. The sets $\{w_{1},v_{0},v_{1}\}$, $\{w_{1},v_{0},v_{2}\}$, $\{w_{1},v_{2},u^{\prime}\}$, and $\{u_{2},v_{0},v_{2}\}$ each induce a triangle. Therefore, $G$ contains $H_{a}$ as a subgraph, a contradiction.
@@
Acknowledgements
Y. Zhang was partially supported by the National Natural Science Foundation of China (NSFC) under Grant No. 11601527 and by the Natural Science Foundation of Hebei Province under Grant No. A2023205045.

### v2

@@
Abstract
Given two graphs $G_{1}$ and $G_{2}$, the Ramsey number $R(G_{1},G_{2})$ denotes the smallest integer $N$ such that any red-blue coloring of the edges of $K_{N}$ contains either a red $G_{1}$ or a blue $G_{2}$. Let $G_{1}$ be a graph with chromatic number $\chi$ and chromatic surplus $s$, and let $G_{2}$ be a connected graph with $n$ vertices. The graph $G_{2}$ is said to be Ramsey-good for the graph $G_{1}$ (or simply $G_{1}$-good) if, for $n\geq s$,
@@ Abstract
keywords:
Ramsey goodness, the Hajós graph, fan
@@ 1 Introduction
The study of Ramsey numbers for sparse graphs has flourished since the 1970s. At that time, Burr (1981) established a general lower bound for the Ramsey number of any pair of graphs. Suppose $G_{2}$ is a connected graph, and let $\chi(G_{1})$ and $s(G_{1})$ denote the chromatic number and the chromatic surplus of $G_{1}$, respectively, where the chromatic surplus refers to the smallest size of a color class over all proper $\chi(G_{1})$-colorings of $G_{1}$. Then, if $|G_{2}|\geq s(G_{1})$,
@@
When equality holds in (1), the graph $G_{2}$ is said to be $G_{1}$-good, or equivalently, $G_{2}$ is Ramsey-good for $G_{1}$. In the special case where $G_{1}$ is the complete graph $K_{k}$, such graphs $G_{2}$ are referred to as $k$-good by Burr and Erdős (1983).
When $G_{1}$ is the complete $(k+1)$-partite graph $K_{1,m_{1},m_{2},\ldots,m_{k}}$ and $G_{2}$ is a star, Burr et al. (1983) proved that for $m=\min\{m_{i}\mid i\in[k]\}$ and sufficiently large $n$,
@@
A frequently studied class of star-like graphs is the fan. A fan $F_{n}$ is the graph formed by $n$ triangles sharing a common vertex; see Figure 1. The concept of the fan was first introduced into extremal graph theory by Erdős et al. (1995), who studied its extremal graph and Turán number. For results on the Ramsey numbers of fans, see Chen et al. (2021); Dvořák and Metrebian (2023); Lin and Li (2009); Zhang et al. (2015).
@@
Li and Rousseau (1996) investigated the Ramsey-goodness of generalized fans. For any graphs $H$ and $G$, they proved that $K_{1}+nH$ is Ramsey-good for $K_{2}+G$ when $n$ is sufficiently large.
Observe that in any proper coloring of the graph $K_{2}+G$, there are two color classes containing only one vertex each. What happens if we require only one color class to have a single vertex? Let $K_{1}+K_{k}(m)$ denote the complete $(k+1)$-partite graph in which one partite set has size $1$ and each of the remaining $k$ partite sets has size $m$. Lin et al. (2010) showed that for $k\geq 2$, if $m$ is odd or if $n|H|$ is odd, then
@@
Subsequently, without relying on the Erdős-Stone-Simonovits stability theorem, Chung and Lin (2025) determined the exact value of $R(K_{1,m_{1},m_{2},\ldots,m_{k}},K_{1}+nH)$ for large $n$ in all cases. Let $m=\min\{m_{i}\mid i\in[k]\}$. Then
@@
Another frequently studied star-like graph is the book $B_{k,n}$, which is the graph $K_{k}+(n-k)K_{1}$. Equivalently, $B_{k,n}$ can be viewed as the graph obtained by blowing up the center of the star $K_{1,n-k}$ into a clique $K_{k}$, while preserving the adjacency between each vertex of $K_{k}$ and the remaining vertices. For results concerning the Ramsey goodness of $B_{k,n}$, we refer the reader to Fox et al. (2023); Nikiforov and Rousseau (2004); Liu and Li (2025). For results on the Ramsey non-goodness of $B_{k,n}$, see Fan and Lin (2023); Fan et al. (2024); Lin and Liu (2021).
@@
At the end of this section, we introduce some concepts and notation that will be used throughout. We use $[n]$ to denote the set $\{1,2,\ldots,n\}$. For a graph $G$ and a vertex subset $X\subseteq V(G)$, the subgraph induced by $X$ is denoted by $G[X]$, whose vertex set is $X$ and whose edge set consists of all edges in $G$ with both endpoints in $X$. For $u\in V(G)$ and $X\subseteq V(G)$, let $N_{X}(u)$ denote the set of neighbors of $u$ in $X$, and let $|N_{X}(u)|$ denote the number of such neighbors, also written as $d_{X}(u)$. When $X=V(G)$, we write $N_{X}(u)$ and $d_{X}(u)$ simply as $N(u)$ and $d(u)$, respectively. The minimum and maximum degrees of $G$ are denoted by $\delta(G)$ and $\Delta(G)$, respectively. Notation and terminology not explicitly defined follow Bondy and Murty (2008).
@@ 2 Proof of Theorem 1
When $n$ is odd, let $\ell=(n+1)/2$. Consider the graph $(\ell K_{2})+(\ell K_{2})$, which consists of two disjoint copies of a matching with $\ell$ edges, where the edges between the two copies form a complete bipartite graph. It is straightforward to verify that this graph does not contain the Hajós graph as a subgraph, and that the maximum degree in its complement is n1n-1. Therefore, $R(H_{a},K_{1,n})\geq 2n+3$ for odd $n$.
@@
By the classical result of Chvátal (1977), we have $R(K_{3},K_{1,n})=2n+1$. Therefore, the graph $G$ must contain a triangle; without loss of generality, let its vertices be $u_{1},u_{2},u_{3}$. For $1\leq i<j\leq 3$, the minimum degree condition implies
@@
We first consider the case when $n$ is even. Since $\delta(G)\geq n+2$, each vertex $w_{i}$ for $i\in[4]$ has at least n1n-1 neighbors in $V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}$. As $|V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}|=2n-2$, there must exist two vertices, say $w_{1}$ and $w_{2}$, that share a common neighbor in $V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}$, denoted by $w_{5}$. Since $G$ does not contain $K_{5}-e$ as a subgraph, neither $w_{3}$ nor $w_{4}$ is adjacent to $w_{5}$.
Let $X=V(G)\setminus\{w_{1},w_{2},w_{3},w_{4},w_{5}\}$. Then |X|=2n3|X|=2n-3. By the minimum degree condition,
@@
Since |X|=2n3|X|=2n-3, $d_{X}(w_{1})\geq n-2$ and $d_{X}(w_{3})\geq n-1$, the sets $N_{X}(w_{1})$ and $N_{X}(w_{3})$ form a partition of $X$, and we have $|N_{X}(w_{1})|=n-2$ and $|N_{X}(w_{3})|=n-1$. Given that $d_{X}(w_{2})\geq n-2$ and $N_{X}(w_{2})\cap N_{X}(w_{3})=\emptyset$, it follows that $N_{X}(w_{1})=N_{X}(w_{2})$. Therefore,
@@
Since $n$ is even, by the classical result of Harary (1972),
@@
Now consider the case where $n$ is odd. Let $Y=V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}$. Since $\delta(G)\geq n+3$, each vertex $w_{i}$ for $i\in[4]$ has at least $n$ neighbors in $Y$. Given that |Y|=2n1|Y|=2n-1, any two vertices among $w_{1},w_{2},w_{3},w_{4}$ share at least one common neighbor in $Y$. Without loss of generality, assume that $w_{1}$ and $w_{2}$ share a common neighbor $w_{5}$ in $Y$, and $w_{2}$ and $w_{3}$ share a common neighbor $w_{6}$ in $Y$. As $G$ does not contain $K_{5}-e$ as a subgraph, vertices $w_{5}$ and $w_{6}$ must be distinct. Consequently, we obtain a Hajós graph in which the vertex subsets $\{w_{1},w_{2},w_{3}\}$, $\{w_{1},w_{2},w_{5}\}$, $\{w_{2},w_{3},w_{6}\}$, and $\{w_{1},w_{3},w_{4}\}$ each induce a triangle. ∎
3 Proof of Theorem 2
The proof of Theorem 2 relies on the following key lemma, which was established by Zeng et al. (2025).
@@ 3 Proof of Theorem 2
To prove the upper bound, we proceed by contradiction. Assume that a graph $G$ with $4n+2$ vertices contains neither $H_{a}$ as a subgraph nor $F_{n}$ as a subgraph in $\overline{G}$. We will derive contradictions in the following two cases.
@@
Let $u$ be a vertex in $\overline{G}$ with the maximum degree. Denote by $H^{{}^{\prime}}$ the subgraph of $\overline{G}$ induced by the set of vertices adjacent to $u$ in $\overline{G}$, i.e., $H^{{}^{\prime}}=\overline{G}[N_{\overline{G}}(u)]$. Let $M$ be a maximum matching in $H^{{}^{\prime}}$. Then $M$ must contain exactly n1n-1 edges. This is because if $M$ has at least $n$ edges, these $n$ edges together with $u$ would form an $F_{n}$ in $\overline{G}$, a contradiction. Conversely, if $M$ has at most n2n-2 edges, the number of vertices in $H^{{}^{\prime}}$ not covered by $M$ (i.e., the remaining vertices after removing the vertices of $M$) is at least 2n+22(n2)=62n+2-2(n-2)=6. These vertices form a complete graph in $G$, which clearly contains $H_{a}$ as a subgraph, leading to a contradiction.
@@
We only need to prove that $u_{1}u_{3}\not\in E(G)$; the other case can be proven similarly. Suppose $u_{1}u_{3}\in E(G)$. Since $\delta(G)\geq 2n$, the set $\{u_{1},u_{2},u_{3}\}$ must contain two vertices, say $u_{1}$ and $u_{3}$, that have a common neighbor in $V(G)\setminus V(W_{4})$. Otherwise, this would contradict the fact that $G$ has $4n+2$ vertices. Let $u_{5}$ be a common neighbor of $u_{1}$ and $u_{3}$ in $V(G)\setminus V(W_{4})$. Then the subgraph induced by the vertex set $\{u_{0},u_{1},u_{2},u_{3},u_{4},u_{5}\}$ in $G$ contains $H_{a}$ as a subgraph, a contradiction. ∎
If $u_{1}$ and $u_{2}$ have a common neighbor distinct from $u_{0}$, denoted by $u_{5}$, then the subgraph induced by the vertex set $\{u_{0},u_{1},u_{2},u_{3},u_{4},u_{5}\}$ contains $H_{a}$ as a subgraph. Therefore, the only common neighbor of $u_{1}$ and $u_{2}$ can be $u_{0}$. Furthermore, since $\delta(G)\geq 2n$, $u_{2}$ must have at least 2n12n-1 neighbors in $V(G)\setminus N(u_{1})$. Similarly, $u_{4}$ must also have at least 2n12n-1 neighbors in $V(G)\setminus N(u_{1})$. Hence, the number of common neighbors of $u_{2}$ and $u_{4}$ in $V(G)\setminus N(u_{1})$ is at least
$2(2n-1)-|V(G)\setminus N(u_{1})|\geq 2(2n-1)-(4n+2-2n)=2n-4.$
@@
Next, we only need to prove that the graph $G[U_{1}\setminus\{u_{1},u_{3}\}]$ does not contain a triangle as a subgraph. Again, we proceed by contradiction. Suppose $G[U_{1}\setminus\{u_{1},u_{3}\}]$ contains a triangle with vertices $x_{1},x_{2},x_{3}$. There must exist two vertices among $\{x_{1},x_{2},x_{3}\}$, say $x_{1}$ and $x_{3}$, that have a common neighbor in $V(G)\setminus\{x_{1},x_{2},x_{3},u_{2},u_{4}\}$, otherwise this would contradict the fact that the number of vertices in $G$ is $4n+2$. Let $x_{4}$ be a common neighbor of $x_{1}$ and $x_{3}$ in $V(G)\setminus\{x_{1},x_{2},x_{3},u_{2},u_{4}\}$. Then the subgraph induced by the vertex set $\{x_{1},x_{2},x_{3},x_{4},u_{2},u_{4}\}$ in $G$ contains $H_{a}$ as a subgraph, which leads to a contradiction. ∎
@@
When $k\geq 2$, for each $1\leq i\leq k-1$, the center vertex of $C_{i}$ and a leaf vertex of $C_{i+1}$ form an edge in $\overline{G}$. Similarly, the center vertex of $C_{k}$ and a leaf vertex of $C_{1}$ form an edge in $\overline{G}$. These $k$ edges form a matching in $\overline{G}$. The remaining vertices in $\overline{G}[U_{1}\setminus\{u_{1}\}]$ induce a complete subgraph, allowing us to find a matching of $n$ edges in $\overline{G}[U_{1}\setminus\{u_{1}\}]$. This gives a fan $F_{n}$ centered at $u_{1}$ in $\overline{G}$, a contradiction.
@@
Next, consider $w_{1},w_{2}\in W_{3}$. By the definition of $W_{3}$, $w_{1}$ and $w_{2}$ each have at least 2n92n-9 neighbors in $U_{2}$. Thus, $w_{1}$ and $w_{2}$ share at least $2(2n-9)-|U_{2}|\geq 2n-18$ common neighbors in $U_{2}$, where this inequality follows from Claim 3.
If each common neighbor of $w_{1}$ and $w_{2}$ in $U_{2}$ has degree at least $2$ in $G[U_{2}]$, note that each connected component in $G[U_{2}]$ has at most one vertex of degree at least $2$. Hence, $G[U_{2}]$ must have at least 2n182n-18 components, each containing a vertex of degree at least $2$, contradicting $|U_{2}|\leq 2n$. Thus, there exists a common neighbor $u_{z}$ of $w_{1}$ and $w_{2}$ in $U_{2}$ with degree at most $1$ in $G[U_{2}]$. Then the vertex $u_{z}$ has at least $2n-2-|W|\geq 2n-11$ neighbors in $U_{1}$. Consequently, $u_{z}$, $w_{1}$, and $w_{2}$ share at least $(2n-11)+(|U_{1}|-5)-|U_{1}|\geq 2n-16$ common neighbors in $U_{1}$. Selecting three such common neighbors together with $u_{z},w_{1},w_{2}$ induces a complete multipartite graph $K_{1,1,1,3}$, which contains $H_{a}$ as a subgraph, a contradiction.
If $w_{1}$ and $w_{2}$ belong to $W_{1}$ and $W_{3}$, respectively, let $w_{1}\in W_{1}$ and $w_{2}\in W_{3}$. Then $w_{1}$ must be adjacent to either $u_{2}$ or $u_{4}$. By symmetry, assume $w_{1}u_{2}\in E(G)$. According to the selection of $W_{3}$, $w_{2}$ has at least 2n92n-9 neighbors in $U_{2}$. Therefore, $w_{2}$ must have a neighbor $u_{z}$ in $U_{2}$ whose degree in $G[U_{2}]$ is at most 1; otherwise, this would contradict $|U_{2}|\leq 2n$. Consequently, the number of neighbors of $u_{z}$ in $U_{1}$ is at least $2n-2-|W|\geq 2n-11$. Thus, the number of common neighbors of $u_{z}$, $w_{1}$, and $w_{2}$ in $U_{1}$ is at least $(2n-11)+(|U_{1}|-5)-|U_{1}|\geq 2n-16$. Selecting two common neighbors, together with $u_{z}$, $w_{1}$, $w_{2}$, and $u_{2}$, forms a subgraph in $G$ that contains $H_{a}$ as a subgraph, leading to a contradiction. ∎
@@
If exactly one vertex in $\{w_{1},\ldots,w_{t}\}$ is not in $V(M_{1})$, assume without loss of generality that $w_{1}\notin V(M_{1})$. By the choice of $M_{1}$, $w_{1}$ is adjacent to all vertices in $U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))$, including $v_{0}$, $v_{1}$, and $v_{2}$. This implies that $w_{1}$ has at least 2n112n-11 neighbors in $U_{1}$.
If $w_{1}\in W_{3}$, by the definition of $W_{3}$, $w_{1}$ has at least 2n112n-11 neighbors in $U_{2}$. Since $v_{2}$ has exactly one neighbor in $U_{1}$, it must have at least $2n-2-|W|\geq 2n-11$ neighbors in $U_{2}$. Thus, $w_{1}$ and $v_{2}$ share at least $2(2n-11)-|U_{2}|\geq 2n-22$ common neighbors in $U_{2}$. Let $u^{\prime}$ be one such common neighbor. The sets $\{w_{1},v_{0},v_{1}\}$, $\{w_{1},v_{0},v_{2}\}$, $\{w_{1},v_{2},u^{\prime}\}$, and $\{u_{2},v_{0},v_{2}\}$ each induce a triangle. Therefore, $G$ contains $H_{a}$ as a subgraph, a contradiction.
@@
Acknowledgements.
The authors are grateful to the anonymous referee for helpful remarks. Y. Zhang was partially supported by the National Natural Science Foundation of China (NSFC) under Grant No. 11601527 and by the Natural Science Foundation of Hebei Province under Grant No. A2023205045.